A particular first-order reaction has a rate constant of 1.35 x 10^2s-1 at 25.0 degrees celsius. What is the magnitude of k at 65.0 degrees celsius if Ea = 55.5 kJ/mol?
Answer #1

Arrehenius equation- ln (K2/k1) = ( -Ea/R) (1/T2- 1/T1)

ln(K2/ 1.35^2) = (-55.5^3/8.314) (1/338 -1/298)

K2= 1.912^2 s-1